OpenStudy (anonymous):

length of a rectangle is 20 meter and its are is 320 squaremeters .find the area of the square drawn on the diagonal of a square whose perimeter equals that of the given rectangle.

OpenStudy (anonymous):

need help..

OpenStudy (anonymous):

still here?

OpenStudy (anonymous):

yups

OpenStudy (anonymous):

You need to find z in this set up given the info. you have.

OpenStudy (anonymous):

You can find z by finding x, and you find x using the information you have about the rectangle.

OpenStudy (anonymous):

The perimeter of the big square is equal to the perimeter of the rectangle.

OpenStudy (anonymous):

If the rectangle has area 320m^2 and one side 20m, then the other side of the rectangle must be found from the formula for the area:\[A=lw \rightarrow 320=20w \rightarrow w = 16\]

OpenStudy (anonymous):

So the perimeter of the rectangle is 16x2 + 20x2 = 72. But we're told this is the perimeter of the big square, so the x-value in the diagram is \[x=\frac{72}{4}=18\]since each side of a square is equal.

OpenStudy (anonymous):

Now we have to use information about the side length (x) and the angles that have been made by inserting that smaller square alongside the diagonal.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

The diagonal of the square bisects the angle it cuts through, so we have the following (see angles):

OpenStudy (anonymous):

I've called one side of one triangle h and the other k, so that x=h+k.

OpenStudy (anonymous):

hmm..ya ok

OpenStudy (anonymous):

In the bottom left triangle, we have\[\sin(45^o)=\frac{z}{h}\]but\[\sin(45^o)=\frac{1}{\sqrt{2}}\]so\[\frac{z}{h}=\frac{1}{\sqrt{2}}\]Also, in the bottom right, you have\[\sin(45^o)=\frac{z}{k}\rightarrow \frac{1}{\sqrt{2}}=\frac{z}{k}\]

OpenStudy (anonymous):

Since sin(45)=z/h=z/k, we must have h=k, and so h+k=2h=18, which means h=9.

OpenStudy (anonymous):

So from any one of those two sine equations, you have\[\frac{z}{h}=\frac{1}{\sqrt{2}}\rightarrow z=\frac{h}{\sqrt{2}}=\frac{9}{\sqrt{2}}\]

OpenStudy (anonymous):

The area of the square is\[z^2=\frac{81}{2}\]

OpenStudy (anonymous):

but lokisan i dont know trignometry...i am in standard 9..

OpenStudy (anonymous):

Hmmmm...let me see if there's another way.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Are you taught Pythagoras' Theorem? I don't know how the Indian system works.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What are you currently learning? That might help me understand what your teacher's thinking.

OpenStudy (anonymous):

mensuration

OpenStudy (anonymous):

Are you 'allowed' to know that the diagonals of a square bisect the right angles?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

man...

OpenStudy (anonymous):

oh yes.. i understood

OpenStudy (anonymous):

I know you understood...I'm just trying to stick to what you're allowed to know.

OpenStudy (anonymous):

understanding has no restrictions ;) ,

OpenStudy (anonymous):

yeah...have you covered congruence tests for triangles?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

Okay, I'll try and put something together with that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I might have something, but I want to double check.

OpenStudy (anonymous):

hey i think i'hv got the answer ..it's coming to 648..is it correct.. ??

OpenStudy (anonymous):

You have to scrap what I wrote before too - I stuffed one of the sines up. And your 648 is a number I'm getting too.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

You can do it using congruent triangles. Only catch is that you need to know that the diagonal of a square bisects the right-angle in to 45 degrees.

OpenStudy (anonymous):

okay.. thanks a lot .need to go and practice some more problems.will bug u if in need.thanx again

OpenStudy (anonymous):

Oh...okay...anyway, I got \[z= \frac{\sqrt{648}}{3}\]so that the area is \[z^2=\frac{648}{9}=72\]

OpenStudy (anonymous):

Congruent triangles and Pythagoras' Theorem were used.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool...well, if you're happy, I'm happy :)